A ray of light is incident at 60degrees in air on air-glass plane surface. Find the angle of refraction in the glass (n for glass = 1.5)


1
Expert's answer
2021-03-08T08:33:19-0500

Let's write the Snell's Law:


"n_isin\theta_i=n_rsin\theta_r,""\theta_r=sin^{-1}(\dfrac{n_isin\theta_i}{n_r}),""\theta_r=sin^{-1}(\dfrac{1.0\cdot sin60^{\circ}}{1.5})=35.3^{\circ}."